A Web music store offers two versions of a popular song. The size of the standard version is 2.9 megabytes (MB). The size of the high-quality version is 4.4 MB. Yesterday, there were 1010 downloads of the song, for a total download size of 3424 MB. How many downloads of the high-quality version were there?
* I tried this problem several times and could not get the answer. Need help!! Thanks :)

Answer by stanbon(75887)

A Web music store offers two versions of a popular song. The size of the standard version is 2.9 megabytes (MB). The size of the high-quality version is 4.4 MB. Yesterday, there were 1010 downloads of the song, for a total download size of 3424 MB. How many downloads of the high-quality version were there?
-------
Downloads Equation: s + h = 1010
Megabyte Equation::2.9s + 4.4h = 3424
-----------------------------
Modify for elimination:
29s + 29h = 29*1010
29x + 44h = 34240
-----------------
Solve for "h":
15h = 4950
h = 330 (# of high-quality downloads)
------------
Cheers,
Stan H.
================